OpenStudy (anonymous):

How can I determine the axis of symmetry, the domain, and range without graphing y = −3x2 − 3x + 4 ?

OpenStudy (anonymous):

comlete the square

OpenStudy (anonymous):

y-4 -3 = −3(x2 + x + 1)

OpenStudy (anonymous):

(y-7) =-3(x+1)^2

OpenStudy (anonymous):

line of symmetry is at x=-1

OpenStudy (anonymous):

thank you ^.^

OpenStudy (anonymous):

you are welcome.